# Examples of robotic art Encountering Robotic Art: The Social, Material, and Temporal Processes of Creation with Machines

### v1

@@ 5. Discussion
This study of robotic art practice provides an empirical description of human-machine relationships in real-world creative processes. This paper first contributes as an empirical description of artistic creativity in robotic art practice—an unconventional use of robots—examined through the artists’ perspectives on their creative experiences. Our analysis reveals three facets of creativity in robotic art practices: the social, material, and temporal. Creativity emerges from the co-constitution between artists, robots, and the audience in spatial-temporal dimensions, as illustrated in Figure 2. Acknowledging the audience as an important actor reflects the social dimension in that creativity does not stem from the artists but from their interactions with the audience. Robots are the major material and technological actants characterizing the practice, and conditioning while illuminating creativity through materiality. The axis of the temporal process signifies that the practice is situated within a time continuum, and the actors/actants with their relations may shift over time. The temporal process thus constitutes another dimension of creativity in robotic art.
Accordingly, as the second contribution of this study, we propose several implications for socially informed, material-attentive, and process-oriented creation with computing systems222For the sake of clarity, we mean “creation with computing systems” by three types of scenarios: human creator(s) create computing system(s) as the final artifact(s) (e.g., robots are artworks themselves); human creator(s) use computing system(s) to create the artifact(s) (e.g., robots create artworks as human planned); and human creator(s) and system(s) simultaneously collaborate during producing the artifact(s) (co-creation). to facilitate creation practices, which can be informative to topics in HCI such as media/art creation, crafting, digital fabrication, and tangible computing. In each following subsection, we explain each implication with a grounding in corresponding findings from our study and relevant literature in HCI and adjacent fields on art, creativity, and creation.
@@ 5.1. Socially Informed Creation
The sociality of creativity means that creativity is distributed among different human actors, commonly within the creators or between the creators and the audience. Glaveanu (2014) found in his ethnographic study on Easter egg decoration in northern Romania, where the artisans would consider how others might appreciate their work and change their creative decisions. The audience can be “present” during the creation process because the audience’s potential appreciation is internalized by the creators, so the creators co-create with the audience through feedback. In this sense, the sociologist Katherine Giuffre asserts that “creative individuals are embedded within specific network contexts so that creativity itself, rather than being an individual personality characteristic is, instead, a collective phenomenon” (Giuffre et al., 2012, p. 1).
@@
We found that the practice of robotic art manifests this sociality as it involves, particularly artists and audiences. Our artists take audiences’ reactions to their artwork as feedback and then revise the robots’ functions and aesthetics accordingly. For example, as shown earlier, Robert added a protective fuse onto his robot because he expected that children would squeeze the springs together and cause a short circuit; Alex’s enthusiasm and attention to the audience’s imagination about his robots led him to new aesthetic designs of both the robots and the scene layouts. The artists may directly ask about the audience’s judgment of quality but they often receive feedback just by observing the audience’s reactions or sometimes by learning from the audience’s imagination about the robots. Meanwhile, our findings reveal that audience reception is not an individual matter but is often associated with their sociocultural codes, including shared cultural norms, beliefs, expectations, and aesthetic values. The audience can be seen as representatives of these broader cultural codes. For example, Mark and Robert felt that the animist cultural tendency of some East Asian societies is associated with higher acceptance of and interest among the audience in intelligence and agency of robots and non-human entities. The sociocultural backgrounds shape the way the audience perceives the work. Such an understanding of expectations informed the artists’ creation.
@@
A creative process, including creation and reception, is essentially a social activity. The second wave of creativity research in psychology has argued for creativity’s dependency on sociocultural settings and group dynamics (Sawyer and Henriksen, 2024). Recent discussions from creativity-support and social computing researchers also called for more attention to the social aspect of creativity (Kato et al., 2023; Fischer et al., 2005, 2009). There is a clear need to consider the audience when producing creative content. For instance, some video-creation-support researchers examined audience preferences of videos to inform system designs to fit the preferences (Wang et al., 2024). These works point to the direction that creative activities involve beyond individual creators to co-creators and heterogeneous audiences. Some HCI researchers conceptualize creativity as by large a socially constructed concept, perceived and determined by social groups (Fischer et al., 2009). Prior HCI work examined the social aspects between art creators. For example, creators and performers in music and dance form social relationships through artifacts, making the final work a collaborative outcome (Hsueh et al., 2019a). There is also a system designed to support collaborative creation between artists (Striner et al., 2022). However, the social creative process between creators and audience is less articulated in HCI. Jeon et al.’s work (Jeon et al., 2019) stands as an exception, suggesting that professional interactive art can involve evaluation with the audience in the creation stage. Another relevant approach in HCI involves enabling the general public to participate in co-creation alongside professional creators. Matarasso (2019), for instance, promoted “participatory art” as “the creation of an artwork by professional artists and non-professional artists working together” with non-professional artists referring to the general public engaged in the art making process. Similarly, socially inclusive community-based art also considers target communities’ perception of the artwork during creation (Clarke et al., 2016, 2014). But like participatory design (Schuler and Namioka, 1993), these art projects aim for social justice more than creativity in the work (Murray et al., 2024), let alone that direct participation in art creation is not always feasible. Our findings suggest that feedback from the audience can lead to creative ideas, as well as that the feedback can be generative and remain low-effort for the audience.
@@
Unlike conventional design feedback, which designers expect to be specific, justified, and actionable (Yen et al., 2024; Krishna Kumaran et al., 2021), good feedback for our artists can be implicit, creative, and generative. Such feedback may include audiences’ imaginations stimulated by the work, future extension ideas, personal and societal reflections, and even emotions, facial expressions, micro-actions, and post-experience behaviors. Our artists gathered this implicit feedback not by asking evaluative questions, as commonly done in typical design processes (e.g., usability testing, think-aloud protocols), which seek to elicit clear, structured responses. Instead, they observe the audience’s reactions and interpret their subjective perceptions. Implicit feedback, while indirect, can lead to more creative ideas than standardized, explicit feedback, as implicity embraces open interpretations of the system (Sengers and Gaver, 2006). Computing systems for creation should better incorporate implicit feedback in addition to explicit ones from the audience into the creation process. Implicit feedback can be indirect, creative, inspirational, and heuristic about functions and aesthetics. A hypothetical instance of such design can be a system that helps creators perceive audiences’ implicit reactions and perceptions and variously interpret them, for further iteration.
@@
Moreover, as seen in Robert and Daniel’s experiences, the audience may participate in robotic live performances by interacting with the robots, who may change actions accordingly, triggering a loop of simultaneous mutual influence that makes the work performative and improvised. HCI researchers explored performative and improvisational creation with machines, focusing on developing and evaluating systems with performative capabilities, including music improvisation with robots (Hoffman and Weinberg, 2010), dance with virtual agents (Jacob and Magerko, 2015; Triebus and Geiger, 2023), and narrative theatre (Magerko et al., 2011; Piplica et al., 2012). Kang et al. (2018) discussed the improvisational nature of interactions between humans and computers and argued that an HCI researcher-designer’s improvisation with the environment facilitates the emergence of creativity and knowledge. Designs of computing systems for creation can leverage performativity for creative experience. One possible direction would be to allow the audience to embed themselves in and interact with elements of static artwork in a virtual space, turning the exhibition into an improvisational on-site creation (Zhou et al., 2023). While interactions with machines during performance are mostly physical or embodied, we posit that they can also be a symbolic engagement. Alex’s audience projected themselves and their personalities onto his robots, which established a symbolic relevance, generating creative imaginations. During exhibitions, East Asian audiences carried their animist views shaped by their sociocultural backgrounds, and robots through the performance were successful in symbolically matching the animist views, stimulating aesthetic satisfaction. Symbolic engagement resonates with what Nam and Nitsche (2014) called the “reference” of the interactive installation performance to participants’ sociocultural conditions. As such, we propose that designers of computing systems for creation may consider establishing symbolic engagement between the produced artifacts and the audience as a way to enhance perceived creativity or enrich the creative experience. One example is an interactive installation, Boundary Functions (Snibbe, 1998), which encourages viewers to reflect on their personal spaces while interacting with the installation and others. Another example is Blendie, a voice-controlled blender that requires a user to “speak” the machine’s language to use it and thus build a symbolic connection with it and make a novel experience (Dobson, 2004).
@@ 5.2. Material-Attentive Creation
The theory of distributed creativity by Glaveanu claims that creativity distributes across humans and materials, so the creation practice itself is inevitably shaped by objects (Glaveanu, 2014). In his case of Easter egg decoration, materials are not passive objects but active participants in artistic creation; e.g., the egg decorators face challenges from color pigments not matching the shell, wax not melted at the desired temperature, to eggs that break at the last step of decoration; hence, materials often go against the decorators’ intentions and influence future creative pathways (Glaveanu, 2014). Materials manifest properties when used for activities; these properties allow some possible uses of the materials while prohibiting others (Leonardi, 2012). Such materiality necessarily shapes how artists create their work. Our findings indicate the critical role of materiality in the artistic practice, based on which we argue that intentional arrangement of materials based on materiality enables the artists to enrich the creative values of their work.
@@
The practice of robotic art depends on specific material properties of robots and other objects. One of the apparent properties of most materials is physicality (Leonardi, 2012), which refers to the fact that materials are physical presences that may interact with other physically present objects. Here, we consider physicality and embodiment interchangeable as computational creativity researchers have conceptualized (Guckelsberger et al., 2021). Our findings support both the conceptual and operational contributions of embodiment for creative activities. For the conceptual aspect, the embodied presence of robotic systems supports creative thinking for our artists, exemplary in Linda’s case where she found new art ideas around the difference between human and robot bodies through bodily engagement with robots. For the operational aspect, the embodied nature of robotic artworks and their creation processes exhibit original aesthetics that are based on physics so different from disembodied works, e.g., embodied drawings by David’s non-industrial robotic arms are dynamic due to physical movements and thus artistically pleasant, which is hard to replicate in simulated programs.
@@
These findings on embodiment of robotic art (Section 4.1) closely relate to HCI’s attention on embodied interaction as a way to leverage human bodies and environmental objects to expand disembodied user experiences. For example, as Hollan et al. (2000) explained, a blind person’s cane or a cell biologist’s microscope as an embodied material is part of the distributed system of cognitive control, showing that cognition is distributed and embodied. Similarly, theories of embodied interaction in HCI explicate how bodily interactions shape perception, experience, and cognition (Marshall et al., 2013; Antle et al., 2011, 2009), backed up by the framework of 4E cognition (embodied, embedded, enactive, and extended) (Wheeler, 2005; Newen et al., 2018). Prior works suggest that creative activities with interactive machines rely on similar embodied cognitive mechanisms (Guckelsberger et al., 2021; Malinin, 2019), which are operationalized by tangible computing (Hornecker, 2011). As related to robots in creation, HCI researchers show that physicality or embodiment of robots in creation may lead to some beneficial outcomes, such as curiosity from the audience, feelings of co-presence, body engagement, and mutuality, which are hard to simulate through computer programs (Dell’Ariccia et al., 2022; Hoggenmueller and Hespanhol, 2020). Embodied robotic motions convey emotional expressions and social cues that potentially enrich and facilitate creation activities like drawings (Ariccia et al., 2022; Grinberg et al., 2023; Dietz et al., 2017; Santos and Egerstedt, 2021). Guckelsberger et al. (Guckelsberger et al., 2021) showed in their review that embodiment-related constraints (e.g., the physical limitations of a moving robotic arm) also can stimulate creativity. These constraints push creators to develop new and useful movements, echoing the broader principle that encountering obstacles in forms or materials can lead to generative processes. This phenomenon is similarly observed in activities such as art and digital fabrication (Devendorf and Ryokai, 2015; Hirsch et al., 2023). In co-drawing with robots, physical touch and textures of drawing materials made the artists prefer tangible mediums (e.g., pencils) than digital tools (e.g., tablets) that fall short in these respects (Jansen and Sklar, 2021).
@@
Materiality plays a crucial role in the embodiment of robots, as the choice of materials fundamentally shapes the physical form and properties. This focus on materials extends to art practices, where robots made with soft materials introduce new aesthetics and sensory experiences (Jørgensen, 2019; Belling and Buzzo, 2021), and the use of plants and soil in robotic printing creates unique visual effects (Harmon et al., 2022). Following Leonardi’s (Leonardi, 2012) conceptualization of materiality, we refer to the materials of robots as encompassing all physical and digital components: shell, hardware, mechanical parts, software, programs, data, and controllers—every element significant to the artist’s intent. Nam et al. (2023) found that the material constraints of robots can limit creative expression but simultaneously stimulate creativity when artists push the boundaries.
@@
Even carefully designed, computational and mechanical components in robots are prone to errors or bugs in everyday runs, causing malfunctions or unexpected consequences. This reflects the unique materiality of robots as complex computing systems. From an engineering perspective, errors signal unreliability and should be eradicated, which is characteristic of advances in robotics focused on error detection and recovery (Gini and Smith, 1987) as well as digital fabrication scenarios that prioritize precision over creativity (Yildirim et al., 2020). However, material failures and accidents are inevitable, exemplifying what has been called the ‘craftsmanship of risk’ (Glaveanu, 2014) in material art. For our artists, these risks are often creatively utilized and incorporated into their work: these moments of breakdown—whether physical or digital—become resources for new creative expression. Errors are anticipated and intentionally designed into the process and work of our artists. In some cases, such as for Alex, the entire concept of one of his works is machine errors.
@@
Thus, to focus on material-attentive creation—considering the creative arrangement of materials—we suggest exploring the embodiment and materiality of creation materials, objects, and environments to recognize their creative potential. This perspective aligns with insights from professional digital fabrication practitioners, who advocate for systems that integrate support for machine settings and material properties alongside geometry and computer simulations (Hirsch et al., 2023). Specifically, we propose using a design method/probe that enables creators to realize both the conceptual and operational contributions of materiality. This approach may build on the material probe developed by Jung and Stolterman (2010), which calls for exploring the materiality of digital artifacts. A material-attentive probe would enable creators to engage with diverse materials, objects, and environments through embodied interaction, encouraging them to speculate on material preferences and limitations, and to compare and contrast material qualities—insights that can inform creative decisions. To accommodate, seek, and actively harness the creative potential of errors, we propose embracing failures, glitches, randomness, and malfunctions in computing systems as critical design materials—elements that creators can intentionally control and manipulate. By doing so, we can begin to systematically approach errors. For instance, as part of the design process, we may document how to replicate these errors and changes, allowing creators to explore them further at their discretion. This could include intentionally inducing errors or random changes to influence the creative process or outcomes.

### v2

@@ 5. Discussion
This study on robotic art explores human-machine relationships in creative processes. It first contributes as an empirical description of artistic creativity in robotic art practice—an unconventional use of robots—examined through the artists’ perspectives on their creative experiences. Our analysis reveals three facets of creativity in robotic art practices: the social, material, and temporal. Creativity emerges from the co-constitution between artists, robots, audience, and environment in spatial-temporal dimensions, as illustrated in Figure 2. Acknowledging the audience as an important actor reflects the social dimension in that creativity does not stem from the artists but from their interactions with the audience. Robots are the major material and technological actants characterizing creative practices, shaping the conditions for creativity to emerge. The axis of the temporal process signifies that the practice is situated within a time continuum, and the actors/actants and their relations shift over time. In this way, temporality constitutes another dimension of creativity in robotic art.
Accordingly, as the second contribution, this study outlines implications for socially informed, material-attentive, and process-oriented creation with computing systems111For the sake of clarity, we mean “creation with computing systems” by three types of scenarios: human creator(s) create computing system(s) as the final artifact(s) (e.g., robots are artworks themselves); human creator(s) use computing system(s) to create the artifact(s) (e.g., robots create artworks as human planned); and human creator(s) and system(s) work in tandem to produce the artifact(s) (e.g., human-robot co-creation). to facilitate creation practices. These insights can inform related HCI research on media/art creation, crafting, digital fabrication, and tangible computing. In each following subsection, we present each implication with a grounding in corresponding findings from our study and relevant literature in HCI and adjacent fields on art, creativity, and creation.
@@ 5.1. Socially Informed Creation
The sociality of creativity means that creativity is distributed among different human actors, commonly within the creators or between the creators and the audience. Glăveanu’s ethnographic study on Easter egg decoration in northern Romania (Glaveanu, 2014) showed that artisans anticipate how others might appreciate their work and adjust their creative decisions accordingly. Even in the absence of direct interaction, the audience’s potential responses become part of the creative process, as artisans imagine feedback and predict reactions. In this sense, the sociologist Katherine Giuffre argues that “creative individuals are embedded within specific network contexts so that creativity itself, rather than being an individual personality characteristic is, instead, a collective phenomenon” (Giuffre et al., 2012, p. 1).
@@
We found that the practice of robotic art manifests this sociality as it involves, particularly artists and audiences. Our artists take audiences’ reactions to their artwork as feedback and then revise the robots’ functions and aesthetics accordingly. For example, as shown earlier, Robert added a protective fuse onto his robot because he expected that children would squeeze the springs together and cause a short circuit; Alex’s enthusiasm and attention to the audience’s imagination about his robots led him to new aesthetic designs of both the robots and the scene layouts. The artists may directly ask about the audience’s judgment of quality but they often receive feedback just by observing the audience’s reactions or sometimes by learning from the audience’s imagination about the robots. Meanwhile, our findings reveal that audience reception is not an individual matter but is often associated with their sociocultural codes, including shared cultural norms, beliefs, expectations, and aesthetic values. The audience can be seen as representatives of these broader cultural codes. For example, Mark and Robert observed that the animist tendency in some East Asian societies is associated with higher acceptance of and interest among the audience in intelligence and agency of robots and non-human entities. Such sociocultural contexts influence not only how audiences interpret the work but also how artists anticipate and respond to these perspectives in their creative process.
@@
A creative process, including creation and reception, is essentially a social activity. The second wave of creativity research in psychology has argued for creativity’s dependency on sociocultural settings and group dynamics (Sawyer and Henriksen, 2024). Recent discussions from creativity-support and social computing researchers also called for more attention to the social aspect of creativity (Kato et al., 2023; Fischer et al., 2005, 2009). There is a clear need to consider the audience when producing creative content. For instance, researchers studying video-creation support have examined audience preferences to inform system designs that align with these preferences (Wang et al., 2024). Such work highlights how creative activities extend beyond individual creators to co-creators and heterogeneous audiences. Some HCI researchers conceptualize creativity as by large a socially constructed concept, perceived and determined by social groups (Fischer et al., 2009). Prior HCI work examined the social aspects between art creators. For example, creators and performers in music and dance form social relationships through artifacts, making the final work a collaborative outcome (Hsueh et al., 2019a). There is also a system designed to support collaborative creation between artists (Striner et al., 2022). However, the social creative process between creators and audience is less articulated in HCI. Jeon et al.’s work (Jeon et al., 2019) stands as an exception, suggesting that professional interactive art can involve evaluation with the audience in the creation stage. Another relevant approach in HCI involves enabling the general public to participate in co-creation alongside professional creators. Matarasso (2019), for instance, promoted “participatory art” as “the creation of an artwork by professional artists and non-professional artists working together” with non-professional artists referring to the general public engaged in the art-making process. Similarly, socially inclusive community-based art also considers target communities’ perception of the artwork during creation (Clarke et al., 2016, 2014). But like participatory design (Schuler and Namioka, 1993), these art projects aim for social justice more than creativity in the work (Murray et al., 2024), let alone that direct participation in art creation is not always feasible. Our findings suggest that feedback from the audience can lead to creative ideas, as well as that the feedback can be generative and remain low-effort for the audience.
@@
Unlike conventional design feedback—which is typically expected to be specific, justified, and actionable (Yen et al., 2024; Krishna Kumaran et al., 2021)—the feedback that resonates with our artists is often implicit, creative, and generative. Such feedback may include audiences’ imaginations stimulated by the work, personal and societal reflections, and even emotions, facial expressions, micro-actions, and observable behaviors following the art experience. Our artists gathered this implicit feedback not by posing evaluative questions, as commonly done in typical design processes (e.g., usability testing, think-aloud protocols), which seek to elicit clear, relatively structured responses. Instead, they closely observe the audience’s reactions and interpret their subjective perceptions. This form of implicit feedback, while indirect, can lead to more creative ideas by embracing open, multifaceted interpretations of the work (Sengers and Gaver, 2006). Computing systems for creation should better incorporate implicit feedback in addition to explicit ones from the audience into the creation process. Implicit feedback can be indirect, creative, inspirational, and heuristic about functions and aesthetics. A hypothetical instance of such design can be a system that helps creators perceive audiences’ implicit reactions and perceptions and variously interpret them, for further iteration.
@@
Moreover, as seen in Robert and Daniel’s experiences, the audience may participate in robotic live performances by interacting with the robots, who may change actions accordingly, triggering a loop of simultaneous mutual influence that makes the work performative and improvisational. HCI researchers explored performative and improvisational creation with machines, focusing on developing and evaluating systems with performative capabilities, including music improvisation with robots (Hoffman and Weinberg, 2010), dance with virtual agents (Jacob and Magerko, 2015; Triebus and Geiger, 2023), and narrative theatre (Magerko et al., 2011; Piplica et al., 2012). Kang et al. (2018) discussed the improvisational nature of interactions between humans and computers and argued that an HCI researcher-designers’ improvisation with the environment facilitates the emergence of creativity and knowledge. Designs of computing systems for creation can leverage performativity in service of creative experience. One possible direction could be to allow the audience to embed themselves in and interact with elements of static artwork in a virtual space, turning the exhibition into an improvisational on-site creation (Zhou et al., 2023). While interactions with machines during performance are mostly physical or embodied, we posit that they can also be a symbolic engagement. Alex’s audience projected themselves and their personalities onto his robots, which established a symbolic relevance, generating creative imaginations. During exhibitions, East Asian audiences carried the animist views shaped by their sociocultural backgrounds, and robots, through the performance, were successful in symbolically matching the views, stimulating aesthetic satisfaction. Symbolic engagement resonates with what Nam and Nitsche (2014) called the “reference” of the interactive installation performance to participants’ sociocultural conditions. As such, we propose that designers of computing systems for creation may consider establishing symbolic engagement between the produced artifacts and the audience as a way to enhance perceived creativity or enrich the creative experience. One example is an interactive installation, Boundary Functions (Snibbe, 1998), which encourages viewers to reflect on their personal spaces while interacting with the installation and others. Another example is Blendie, a voice-controlled blender that requires a user to “speak” the machine’s language to use it. This interaction builds a symbolic connection between the user and the device, transforming the act of blending into a novel experience (Dobson, 2004).
@@ 5.2. Material-Attentive Creation
The theory of distributed creativity by Glaveanu claims that creativity distributes across humans and materials, so the creation practice itself is inevitably shaped by objects (Glaveanu, 2014). In his case of Easter egg decoration, materials are not passive objects but active participants in artistic creation; e.g., the egg decorators face challenges from color pigments not matching the shell, wax not melted at the desired temperature, to eggs that break at the last step of decoration; hence, materials often go against the decorators’ intentions and influence future creative pathways (Glaveanu, 2014). Materials manifest specific properties, which afford certain uses of the materials while constraining others (Leonardi, 2012). Our findings highlight the critical role of materiality in artistic practice, showing that artists intentionally arrange materials to enhance the creative values of their work.
@@
Robotic art relies on the material properties of robots and other objects. An apparent property of most materials is their physicality (Leonardi, 2012), meaning they possess a tangible presence that enables interaction with other physical entities. Here, we consider physicality and embodiment interchangeable as computational creativity researchers have conceptualized (Guckelsberger et al., 2021). Our findings support both the conceptual and operational contributions of embodiment for creative activities. For the conceptual aspect, the embodied presence of robotic systems supports creative thinking for our artists, exemplary in Linda’s case where she found new art ideas around the difference between human and robot bodies through bodily engagement with robots. For the operational aspect, the embodied nature of robotic artworks and their creation processes exhibit original aesthetics that are based on physics much different from disembodied works, e.g., embodied drawings by David’s non-industrial robotic arms are dynamic due to physical movements and thus artistically pleasant, which is hard to replicate in simulated programs.
@@
These findings on embodiment of robotic art (Section 4.1) closely relate to HCI’s attention on embodied interaction as a way to leverage human bodies and environmental objects to expand disembodied user experiences. For example, as Hollan et al. (2000) explained, a blind person’s cane and a cell biologist’s microscope as embodied materials are part of the distributed system of cognitive control, showing that cognition is distributed and embodied. Similarly, theories of embodied interaction in HCI explicate how bodily interactions shape perception, experience, and cognition (Marshall et al., 2013; Antle et al., 2011, 2009), backed up by the framework of 4E cognition (embodied, embedded, enactive, and extended) (Wheeler, 2005; Newen et al., 2018). Prior works suggest that creative activities with interactive machines rely on similar embodied cognitive mechanisms (Guckelsberger et al., 2021; Malinin, 2019), which are operationalized by tangible computing (Hornecker, 2011). As related to robots in creation, HCI researchers show that physicality or embodiment of robots in creation may lead to some beneficial outcomes, such as curiosity from the audience, feelings of co-presence, body engagement, and mutuality, which are hard to simulate through computer programs (Dell’Ariccia et al., 2022; Hoggenmueller and Hespanhol, 2020). Embodied robotic motions convey emotional expressions and social cues that potentially enrich and facilitate creation activities like drawings (Ariccia et al., 2022; Grinberg et al., 2023; Dietz et al., 2017; Santos and Egerstedt, 2021). Guckelsberger et al. (Guckelsberger et al., 2021) showed in their review that embodiment-related constraints (e.g., the physical limitations of a moving robotic arm) can also stimulate creativity. These constraints push creators to develop new and useful movements, echoing the broader principle that encountering obstacles in forms or materials can lead to generative processes. This phenomenon is similarly observed in activities such as art and digital fabrication (Devendorf and Ryokai, 2015; Hirsch et al., 2023). In co-drawing with robots, physical touch and textures of drawing materials made the artists prefer tangible mediums (e.g., pencils) than digital tools (e.g., tablets) that fall short in these respects (Jansen and Sklar, 2021).
@@
Materiality plays a crucial role in the embodiment of robots, as the choice of materials fundamentally shapes the physical forms and properties. This focus on materials extends to art practices, where robots made with soft materials introduce new aesthetics and sensory experiences (Jørgensen, 2019; Belling and Buzzo, 2021), and the use of plants and soil in robotic printing creates unique visual effects (Harmon et al., 2022). Following Leonardi’s (Leonardi, 2012) conceptualization of materiality, we refer to the materials of robots as encompassing physical and digital components—including the shell, hardware, mechanical parts, software, programs, data, and controllers—each significant to the artist’s intent. Nam et al. (2023) found that the material constraints of robots can limit creative expression but simultaneously stimulate creativity when artists push the boundaries.
@@
Even carefully designed, digital and mechanical components in robots are prone to errors or bugs in everyday runs, causing malfunctions or unexpected consequences. This reflects the unique materiality of robots as complex computing systems. From an engineering perspective, errors signal unreliability and must be eliminated, driving advancements in robotics—where error detection and recovery are central (Gini and Smith, 1987)—as well as in digital fabrication, which prioritizes precision over creative exploration (Yildirim et al., 2020). However, material failures and accidents are inevitable, exemplifying what has been called the ‘craftsmanship of risk’ (Glaveanu, 2014) in material art. For our artists, these risks are often creatively utilized and incorporated into their work: these moments of breakdown—whether physical or digital—become resources for new creative expression. Errors are anticipated and intentionally designed into the process and work of our artists. In some cases, such as for Alex, the entire concept of one of his works is machine errors.
@@
Thus, to focus on material-attentive creation—considering the creative arrangement of materials—we suggest exploring the embodiment and materiality of creation materials, objects, and environments to recognize their creative potential. Specifically, we propose using a design method/probe that enables creators to realize both the conceptual and operational contributions of materiality. This approach may build on the material probe developed by Jung and Stolterman (2010), which calls for exploring the materiality of digital artifacts. A material-attentive probe would enable creators to engage with diverse materials, objects, and environments through embodied interaction, encouraging them to speculate on material preferences and limitations, and to compare and contrast material qualities—insights that can inform creative decisions. To accommodate, seek, and actively harness the creative potential of errors, we propose embracing failures, glitches, randomness, and malfunctions in computing systems as critical design materials—elements that creators can intentionally control and manipulate. By doing so, we can begin to systematically approach errors. For instance, as part of the design process, we may document how to replicate these errors and changes, allowing creators to explore them further at their discretion. This could include intentionally inducing errors or random changes to influence the creative process or outcomes.